OpenStudy (anonymous):

What is the dimension of space: problem 1: P (subscript 'n') in [-1,1] problem 2: P in [-1,1] problem 3: P (subscript '3') in [-1,1]

OpenStudy (anonymous):

It's hard to tell exactly since you haven't defined what P is. I'm assuming it's a polynomial? I'm also thinking the subscript is the big giveaway re. dimension. \[\dim(P_n)=n\]\[\dim(P)=1\]\[\dim(P_3)=3\]...if I had to take an educated guess.

OpenStudy (anonymous):

hi Lokisan, thanks for the reply...I was going through the following link http://tutorial.math.lamar.edu/Classes/LinAlg/Basis.aspx and my guess is that dim(Pn)=n+1 dim(P)=1 and dim(P3)=4 I dont know even my prof did not define what p is...may be I have to go around basis and linear independence and try to get the dimension....can u tell me if am thinking in the right direction...

OpenStudy (anonymous):

Sorry - I did mean to put dim(P_n)=n+1 etc.

OpenStudy (anonymous):

It's late.

OpenStudy (anonymous):

so lokisan...is my answer correct...

OpenStudy (anonymous):

If they are polynomials, then it's the case\[\dim(P_n)=n+1\]

OpenStudy (anonymous):

oh ok...thanks a lot...u have any online link...can u share it with me...

OpenStudy (anonymous):

It will depend on the convention for subscripting on the polynomial.

OpenStudy (anonymous):

hmm ok...

OpenStudy (anonymous):

Regarding links, I don't have any atm. All I have are books :)

OpenStudy (anonymous):

I would say,\[\dim(P_n)=n+1\]\[\dim(P_3)=4\]and as for \[\dim(P)\]if you're to take the subscript as 1, then the dimension is 2, and if the subscript is taken to be 0, then the dimension would be 1.

OpenStudy (anonymous):

The link you sent me is pretty good. A lot of people reference those notes.

OpenStudy (anonymous):

hmm ya...that was helpful thanks!

OpenStudy (anonymous):

Is there any subscript on problem 2?

OpenStudy (anonymous):

nope...it was written the same way in my class...may be i put it as how u told...like 1 if the subscript is 0 and 2 if the subscript is 1....i think that shud be fine...isnt it so...

OpenStudy (anonymous):

Oh, okay, if it's a written submission, I would do that. I was just worried in case it was something you had to submit online.

OpenStudy (anonymous):

hmm no...it is a written submission...

OpenStudy (anonymous):

You should probably check with your professor about the convention before submitting. No point losing marks for nothing.

OpenStudy (anonymous):

hmm yeah...

OpenStudy (anonymous):

So, have I answered your question?

OpenStudy (anonymous):

/are you happy with your answers?

OpenStudy (anonymous):

ha ha more than happy! no need to ask...;)

OpenStudy (anonymous):

Good. Just thinking, you might want to check out the MIT lectures on YouTube. They also provide examples which can be helpful. I haven't checked them out for vector spaces, but I think it's worth a try. The examples I've seen on other topics are quite good.

OpenStudy (anonymous):

ya..i checked through Prof.Strang's 18.06 lecture 9....that was helpful too....thanks for ur advice...

OpenStudy (anonymous):

np...good luck :)